OpenStudy (anonymous):

12|1/2y+3|=6 i think its no solution, but i needa be sure...

OpenStudy (anonymous):

hi goodman i love ur pic

OpenStudy (saifoo.khan):

\[|\frac12y+3| = \frac{6}{12}\] \[|\frac12 y + 3| = 0.5\] \[\frac12y + 3 = 0.5 \text{ and }\frac12y + 3 = -0.5 \] Solve both.

OpenStudy (anonymous):

thnks violiin player..urs great to! so wait..i got there too, but i think its no solution..

OpenStudy (anonymous):

thnks

OpenStudy (anonymous):

cuz at skool, a lot of pple got no solutions for tht problem and i dont understand why

OpenStudy (saifoo.khan):

Then it must be correct. :p

OpenStudy (anonymous):

okay..thnks :D

OpenStudy (saifoo.khan):

i think they are wrong let me check.

OpenStudy (saifoo.khan):

And yes they are ..

OpenStudy (saifoo.khan):

Solve it.. wu will get two answers..

OpenStudy (anonymous):

I KNEW IT!! okay..i was jus checking, thnk u

OpenStudy (saifoo.khan):

welcome.